OpenStudy (anonymous):

simplify 9 radical 40

OpenStudy (anonymous):

\(40=4\times 10\) so \(\sqrt{40}=\sqrt{4\times 10}=\sqrt{4}\times \sqrt{10}=2\times \sqrt{10}\) is a start

OpenStudy (anonymous):

thank you :)